OpenStudy (anonymous):

The formula for the volume of a cone is V= 1/3Bh. Find h (in centimeters) if V=176cm^3 and B=40cm^2

OpenStudy (anonymous):

i dont think my answer is going to be right. im looking at it i dont think its the right answer. let me ask my friend here real quick.

OpenStudy (anonymous):

okay thankk

OpenStudy (anonymous):

Alrii so this is all about filling the variables. V = 176cm^3 and B = 40cm^2 so, V = 1/3Bh V(1) = 3Bh V/3B = h h = 176/3(40) h = 1.5 (rounded to the nearest tenth)

OpenStudy (anonymous):

1.5 cm*

OpenStudy (anonymous):

thank you so much @trackzz

OpenStudy (anonymous):

sorry it took me a while, i kept getting mixed up with the rearranging process

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

If you need any other help, im right here :)

OpenStudy (anonymous):

:)) thanks cause agebra2 is killing me

OpenStudy (anonymous):

aha no worries, it won't kill you now :)